Mountaintop golf course is planning for the coming season. Investors would like to earn a​ 12% return on the​ company's

$48,000,000

of assets. The company primarily incurs fixed costs to groom the greens and fairways. Fixed costs are projected to be

$23,000,000

for the golfing season. About

450,000

golfers are expected each year. Variable costs are about

$16

per golfer. Mountaintop golf course is a

priceminus−taker

and​ won't be able to charge more than its competitors who charge

$125

per round of golf. What profit will it earn as a percent of​ assets?

A.Loss of

85.33​%

B.Loss of

54.27​%

C.Profit of

158.94​%

D.Profit of

54.27​%
Answer #1

Correct answer---- (D).Profit of 54.27​%

A

Expected golfers

450000

B

Revenue (450000 x $125)

$ 5,62,50,000.00

C

Variable cost (450000 x 16)

$      72,00,000.00

D=B-C

Contribution margin

$ 4,90,50,000.00

E

Fixed cost

$ 2,30,00,000.00

F=D-E

Profit

$ 2,60,50,000.00

G

Assets

$ 4,80,00,000.00

H=F/G x 100

Return on assets

54.27%